How many gallons of a 3% salt solution must be mixed with 10 gallons of a 7% solution to obtain a 4% solution?
Answer by mananth(16946)

3% solution - x gallons
7% 10 gallons
mixture 4% (x+10)gallons
.....
0.03x+0.07*10=0.04(x+10)
0.03x+0.7=0.04x+0.1
0.04x-0.03x=0.7-0.1
0.01x=0.6
divide by 0.01
x=0.6/0.01
x=6 gallons.